### v3

@@ 7.3 Numerical Studies
We implement our Algorithm 2 to study the numerical performance. We consider a CMDP problem with the state space $|\mathcal{S}|=10$ and the action space $|\mathcal{A}|=10$. We set the discount factor $\gamma=0.7$. We then randomly generate the probability transition kernel $P$. To be specific, for each state $s\in\mathcal{S}$, action $a\in\mathcal{A}$, and the future state $s^{\prime}\in\mathcal{S}$, we uniformly generate a randomly variable $p_{s,a,s^{\prime}}$. Then, the transition probability is defined as $P(s^{\prime}|s,a)=\frac{p_{s,a,s^{\prime}}}{\sum_{s^{\prime\prime}\in\mathcal{S}}p_{s,a,s^{\prime\prime}}}$. For each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected reward $\hat{r}(s,a)$ is uniformly generated from the interval $[1,2]$ (with the reward for the first action set to be 00). The actual reward $r(s,a)=\hat{r}(s,a)+\eta$, where $\eta$ is uniformly distributed among $[-0.5,0.5]$. There are $K=5$ constraints and for each constraint $k\in[K]$ and each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the expected cost $\hat{c}_{k}(s,a)$ to be uniformly generated from $[1,2]$. The actual cost $c_{k}(s,a)=\hat{c}_{k}(s,a)+\eta^{\prime}$, where $\eta^{\prime}$ is uniformly distributed among $[-0.5,0.5]$.

### v4

@@ 9 Numerical Studies
We implement our Algorithm 3 to study the numerical performance. We consider a CMDP problem with the state space $|\mathcal{S}|=10$ and the action space $|\mathcal{A}|=10$. We set the discount factor $\gamma=0.7$. We then randomly generate the probability transition kernel $P$. To be specific, for each state $s\in\mathcal{S}$, action $a\in\mathcal{A}$, and the future state $s^{\prime}\in\mathcal{S}$, we uniformly generate a randomly variable $p_{s,a,s^{\prime}}$. Then, the transition probability is defined as $P(s^{\prime}|s,a)=\frac{p_{s,a,s^{\prime}}}{\sum_{s^{\prime\prime}\in\mathcal{S}}p_{s,a,s^{\prime\prime}}}$. For each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected reward $\hat{r}(s,a)$ is uniformly generated from the interval $[1,2]$ (with the reward for the first action set to be 0). The actual reward $r(s,a)=\hat{r}(s,a)+\eta$, where $\eta$ is uniformly distributed among [0.5,0.5][-0.5,0.5]. There are $K=5$ constraints and for each constraint $k\in[K]$ and each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the expected cost $\hat{c}_{k}(s,a)$ to be uniformly generated from $[1,2]$. The actual cost $c_{k}(s,a)=\hat{c}_{k}(s,a)+\eta^{\prime}$, where $\eta^{\prime}$ is uniformly distributed among [0.5,0.5][-0.5,0.5].